OpenStudy (anonymous):

find the derivative of the function, h(x)=2/(3x)^2

OpenStudy (anonymous):

\[ h(x)=\frac2{(3x)^2}=\frac2{9x^2}=\frac29x^{-2}\\ h'(x)=\frac29(-2x^{-3})=-\frac4{9x^2} \]

OpenStudy (anonymous):

is this a rule>? @oldrin.bataku

OpenStudy (anonymous):

@jessiedeee I used the fact that the derivative of the product of a constant and a function is equivalent to the product of said constant and the derivative of said function, i.e. d/dx [ku] = k d/dx [u]. I then used the power rule, i.e. d/dx x^n = n x^(n-1)

OpenStudy (anonymous):

ohhh ok I see thankyou:)@oldrin.bataku

OpenStudy (anonymous):

thats not rite

OpenStudy (anonymous):

@Dwade03 ?

OpenStudy (anonymous):

we need to use the quotient rule for this problem

OpenStudy (anonymous):

@Dwade03 how would you use the quotient tule for this?

OpenStudy (anonymous):

because u have a function being divided..in this case its the (3x)^2

OpenStudy (anonymous):

ohhh yeahhh! so in that case, f(x)=2 g(x)=(3x)^2 ?? right? @Dwade03

OpenStudy (anonymous):

yes sir